A woman takes her dog Rover for a walk on a leash. To get the little dog moving forward, 

she pulls on the leash with a force of 20.0 N at an angle of 37° above the horizontal. (a) 

How much force is tending to pull Rover forward? (b) How much force is tending to lift 

Rover off the ground? Show graphical and analytic solutions.


1
Expert's answer
2021-03-14T19:13:29-0400

(a) We can find the force that is tending to pull Rover forward (or the horizontal component of the force) as follows:


"F_{forward}=Fcos\theta=20\ N\cdot cos37^{\circ}=16\ N."

(b) We can find the force that is tending to pull Rover off the ground (or the vertical component of the force) as follows:


"F_{up}=Fsin\theta=20\ N\cdot sin37^{\circ}=12\ N."